# Kelvin waves in nonequilibrium universal dynamics of relativistic scalar field theories

### v1

@@ III Excitations and defects in O(1) theory
Figure 1: Left: Phase angle of the complex quantity $\theta=\phi+\mathrm{i}\pi$ defined from the relativistic degrees of freedom. Vortices/antivortices occur at a phase change ±$\pm 2\pi$, as magnified by the middle insets. Right: Vortex defects also show up as distinct minima in the local energy density on the lattice. We show coarse-grained plots to minimise the influence of lattice artefacts. Lengths are shown in units of $Q^{-1}$, cf. Eq. (3). The healing length $\xi_{\mathrm{h}}=p_{\xi_{\text{h}}}^{-1}=[2g\rho_{0}M]^{-1/2}$, which gives the scale of the diameter of the vortex cores, is $\xi_{\text{h}}\approx 0.83\,Q^{-1}$.
@@ III.1 Low-energy effective theories of scalar fields
Relativistic $\mathrm{O}(N)$ scalar field theories are known to display nonrelativistic effects at low momenta in the presence of a mass gap. For the initial conditions considered here, such a mass gap $|p|\lesssim M$ with effective mass $M$ emerges even when the underlying theory is originally massless [28], giving rise to nonrelativistic dynamics in the infrared. Hence, a single component field theory $\mathrm{O}(1)$ can be described by an emergent $\mathrm{U}(1)$ theory. This mapping can be made more rigorous by introducing nonrelativistic degrees of freedom emerging from the underlying relativistic ones
@@
with $\omega_{x}=\sqrt{M^{2}-\nabla^{2}}$ [145, 76]. The emergent dynamics is governed by the Gross-Pitaevskii equation,
@@
thus describing the real-time dynamics of a single-component Bose gas. While it is a well-known mapping, Ref. [76] considered it for the case of overoccupied isolated systems corroborated by numerical simulations in two dimensions, with initial conditions very similar to what is being considered here.
In contrast, for a nonrelativistic $N$-component Bose gas, a low-energy effective theory was considered in Ref. [81], describing interacting Goldstone modes of the total and relative phase excitations. There are $N$ equations of motion for the $\mathrm{U}(N)$ theory, which give rise to $N-1$ Goldstone modes [^1] Based on the underlying $U(N)$ theory, there are $12N-$ 1 broken generators, and $22N-$ 2 of them combine into the $N-1$ quadratic modes [81]. with quadratic dispersion $\omega_{\mathrm{G}}(p)$ (relative phase excitations), and a Bogoliubov mode (total phase) with dispersion $\omega_{\mathrm{B}}(p)$,
@@
where $\rho_{0}$ is the condensate density in the zero-mode. In Ref. [31], these dispersions were generalised to the relativistic case of O($N$) models by computing the dispersions of possible excitations using classical field equations of motion. This gives rise to the gapless modes
@@
where $g\rho_{0}$ imprints the magnitude of the (local) condensate and is related to the effective mass $M$. These dispersion relations are provided as relative excitations on top of the local condensate that is the solution of the classical equation of motion. Therefore, they can be directly compared to the dispersions (16) of the nonrelativistic theories. However, we emphasise that in the relativistic case, one computes correlations of the scalar field $\phi$ instead of $\psi$ and, before comparing to the stated dispersions, one has to remove the rotation frequency ±$\pm M$ in the frequency spectrum [31] that also emerges in the exponent of Eq. (14). We will take this into account in all the figures below by plotting $\omega-M$.
@@
We note that in the large-$N$ limit, the Bogoliubov mode is suppressed in the infrared. In this limit, large-$N$ kinetic theories [28, 29, 30] that dominantly describe the scattering of quasiparticles with a quadratic dispersion agree with the low-energy effective theory (EFT) considered in Ref. [81]. For the single-component scalar field theory, the EFT considers only Bogoliubov modes and no quadratic excitations. However, the extraction of the excitation spectrum in $\mathrm{U}(1)$ and $\mathrm{O}(1)$ theories in Refs. [122, 31] reveals the existence of a dominant “transport peak” or “non-Lorentzian peak” at low momenta, in addition to the Bogoliubov modes, that drives the dynamics. This raises the question of what kind of quasiparticles are present in a single-component scalar field theory and, in particular, what the dominant degree of freedom is that none of the existing EFTs or kinetic theories are able to capture.
@@ III.2 Topological defects in $\mathrm{O}(1)$ scalar field theory
Approaching from a different perspective, an important consideration is the potential existence of topological defects in our systems. This could be especially significant for the $N=1$ case, since the lower $N$ is, the more constrained the topology of the underlying vacuum manifold becomes, making it more likely to support nontrivial topological structures. [^2] For a mathematically more rigorous consideration of what plays the role of a vacuum manifold in the nonequilibrium dynamical theory considered here, see Refs. [61, 102].
Due to the effective conservation of particle number in the infrared regime, which also occurs in the presence of a mass gap in the relativistic theories considered here, an inverse self-similar particle transport develops. This gradually transfers particles to low-momentum modes, ultimately building a condensate at zero momentum. However, causality imposes a limit on the rate at which information and correlations can propagate, preventing the instantaneous formation of a uniform condensate. Therefore, the condensate initially lacks long-range order, which it will gradually develop through ordering dynamics [114]. This dynamics may include topological defects depending on the number of field components $N$ [61], and for a single component, $N=1$, there are stable vortex defects. In this particular case, the formation of vortices and antivortices, and their annihilation process is of importance, since this allows for the growth of larger structures.
One way to identify these defects in our simulations would be performing the canonical transformation (14) on the relativistic degrees of freedom $\phi$ and $\pi$ to obtain the complex $\psi(t,\mathbf{x})$. Then positions on the real-space lattice where depressions in the density of $|\psi(t,\mathbf{x})|^{2}$ occur along with a phase change of approximately ±$\pm 2\pi$ in $\arg[\psi(t,\mathbf{x})]$ would correspond to vortices and antivortices, respectively [76]. However, we find that it is also possible to obtain topological defects directly from the relativistic degrees of freedom $\phi$ and $\pi$ without transforming into $\psi$, and also without applying any cooling effects in the simulations. For a relativistic theory with a sufficiently large mass gap, defects also show up as minima in energy density fluctuations [102], which coincide with a ±$\pm 2\pi$ winding in the relativistic phase angle defined by
@@
as shown in Fig. 1 for the two-dimensional case. We have cross-checked that the number and positions of (anti)vortices coincide in these two observables, the relativistic $\theta(t,\mathbf{x})$, and the nonrelativistic $\arg[\psi(t,\mathbf{x})]$. Such phase singularities in the nonrelativistic $\mathrm{arg}[\psi(t,\mathbf{x})]$ have been specifically matched with depressions in the complex field $\psi$ in Refs. [74, 76] and their time evolution has been tracked in the context of vortex-antivortex annihilation events and the resulting coarsening of the domains. In this work, we will obtain results from purely relativistic quantities, namely the fields $\phi(t,\mathbf{x})$ and $\pi(t,\mathbf{x})$ entering $\theta(t,\mathbf{x})$.
Technically, one could first identify the minima in the energy density and check for a corresponding phase winding in $\theta(t,\mathbf{x})$. In practice, however, nearly all the ±$\pm 2\pi$ phase windings coincide with a minimum in the energy density, such that simply checking only one of these observables gives an accurate picture of the total vortex density. Similarly, for the three-dimensional case, vortex lines are revealed by looking for depressions in the energy density fluctuations, as seen in Fig. 2 for different times.
Figure 3: Initial spatial configurations of the fields $\phi(t_{0},\mathbf{r})$ and $\pi(t_{0},\mathbf{r})$ in $d=2$ dimensions, $\mathbf{r}=(x,y)$, normalised to have values between $-1$ and $1$. The “scrambled” pattern, containing many small domains, allows for numerous phase singularities in $\arg[\phi+\mathrm{i}\pi]$, which we identify as (anti)vortex defects. These domains then become larger, as throughout the coarsening process vortices annihilate with antivortices resulting in the growth of structures. The healing length is $\xi_{\text{h}}\approx 0.83\,Q^{-1}$ in the units shown.
Figure 4: Distribution function $f(t,p)$, cf. (6), for a single-component scalar field theory in (a) two and (b) three dimensions, shown at five different evolution times $t$, where quantities are shown in units of $Q$, i.e., time in units of $Q^{-1}$ and momentum in units of $Q$. The inset shows the functions rescaled according to (22), with $\beta=0.25$ and $\alpha=0.5$ in two dimensions, and $\beta=0.5$ and $\alpha=1.5$ in three dimensions. In (a), simulations were run on a $2048^{2}$ lattice, while in (b) the lattice size was $512^{3}$. The momentum scales $p_{\xi_{\text{h}}}=\sqrt{2g\rho_{0}M}$ corresponding to the healing length $\xi_{\mathrm{h}}=p_{\xi_{\text{h}}}^{-1}$ are $p_{\xi_{\text{h}}}\approx 1.2\,Q$ in $d=2$ and $p_{\xi_{\text{h}}}\approx 0.6\,Q$ in $d=3$.
@@
It is also important to note that using a smoother cutoff for the initial momentum space distribution instead of a step function, like a hyperbolic tangent function or a Gaussian function, did not reduce the number of vortex defects initially formed in the system, which we explicitly checked. As long as one has a relatively steep increase in occupation number in momentum space near some characteristic momentum, this initially scrambled picture will emerge in position space, and vortices will always form very quickly in the dynamics. In this sense, it is not possible not to have, or to remove these defect configurations. For the case of $N=1$ (and also $N=2,3$ but not higher $N$), they are stable configurations on topological grounds [61, 102].
@@ III.3 Excitations on vortices
Another type of contribution yet to be considered is that the vortex defects themselves can support excitations, which appear as quasiparticles during the dynamics. For instance, phonons can be produced from vortex-antivortex annihilation processes [146], which are expected to yield a linear dispersion relation at low momenta. Beyond this, individual vortex lines in three dimensions, and vortex cores in two dimensions can be excited in a way that can be characterised as quasiparticles, which are distinct from the phonons arising from vortex annihilation. In anticipation of our findings, we briefly summarise the properties of such kelvon quasiparticles.
In three dimensional systems, vortex defects are extended objects and may be viewed as vortex lines, along which helical excitations can propagate as waves. These are known as Kelvin waves [124], and were originally described in the context of classical hydrodynamics. In the realm of quantum fluids, Kelvin waves arise as helical perturbations of quantised vortex lines, with the circulation being discretised as a direct consequence of the quantum mechanical phase winding around the vortex core [125, 147, 148, 127]. These excitations play a crucial role in the dynamics of quantum turbulence, mediating energy transfer from larger to smaller scales [128, 129, 17, 130, 131, 132, 133, 19, 134, 135, 20, 136], as Kelvin waves at a similar scale interact, exciting smaller scale Kelvin waves, ultimately dissipated by phonon emission [129, 131, 128]. At larger scales, vortex reconnection is thought to dominate energy transfer [149], and such reconnection events excite Kelvin waves at scales close to the inter-vortex distance. Lord Kelvin obtained, for the small-amplitude wave excitations of a thin columnar vortex with a hollow core, with wave number $p$ along the vortex line and azimuthal quantum number $n=1$, the dispersions [124]
@@
where $\gamma\approx 0.577216\dots$ is the Euler-Mascheroni constant. The Kelvin-wave dispersion (21) also applies to vortices in a superfluid [127] with integer winding number ${w}\in\mathbb{Z}$ and circulation $\Gamma={w}h/m$ quantised in integer multiples of the Planck constant $h$ over particle mass. The much larger $\omega^{+}(p\to 0)\simeq\Gamma/(\pi r_{\mathrm{c}}^{2})$ is related to the Magnus force and the mass inside the core [150].
@@
The dynamics of quantised vortices and their Kelvin-wave excitations with dispersion (20) is well captured by the Gross–Pitaevskii equation (GPE) (15) [151, 152, 138, 153, 154]. As it is explicitly visible in the approximate expression (21), the $p$-dependence of the dispersion (20) is nearly quadratic at low momenta. Around $pr_{\mathrm{c}}=1$ it is close to linear in $p$, while at higher momenta, it bends to a lower power than linear. Such a momentum dependence serves as a unique signature of Kelvin waves that is unlikely to be attributed to any other excitation with only linear or quadratic dispersions. Previous GPE studies of Kelvin waves mostly focused on weak wave turbulence, and a direct connection has not been made with nonthermal fixed points to date, where strong turbulence, including vortex tangles, dominates the infrared region.
@@
At first sight, Kelvin waves cannot occur in two dimensions, where vortices lack a linear extension along the core and are localised objects with circular symmetry. Similar quasiparticles, however, emerge even in two dimensions [139, 155], especially if the vortices are not point-like, but have a finite core size, which is precisely our case, as seen in the right panel of Fig. 1 depicting the energy density. As a result, the boundary of the vortex supports bound-state-like oscillations, which, in the azimuthal direction, are quantised in the same way as kelvons in three dimensions, looked at in a single fixed plane perpendicular to the vortex core. The lowest such azimuthal excitation corresponds to a circularly varying displacement of the centre of the vortex. As a result, also the density is modified in its radial dependence relative to the rotating position of the vortex, and in the large-volume limit, a continuum of kelvon states emerges, which can be distinguished again by a wave number $p$ parametrising the radial dependence as well as the dispersion of the kelvon. Therefore, an analogy can be drawn between a Kelvin wave extended along a vortex line in three dimensions and a “Kelvin wave” representing, at lowest order, a circularly rotating displacement of the vortex position. We will also refer to these quasiparticles in two dimensions as kelvons, for which a dispersion with the same scaling (21) as in three dimensions has been found [139, 155]. Wave-like motion along vortex boundaries has also been discussed in the context of Kelvin waves in finite-size, quasi two-dimensional systems as realised in a trapping potential strongly confined in the third dimension [139, 138].

### v2

@@ III Excitations and defects in O(1) theory
Figure 1: Left: Phase angle of the complex quantity $\theta=\phi+\mathrm{i}\pi$ defined from the relativistic degrees of freedom. Vortices/antivortices occur at a phase change $\pm 2\pi$, as magnified by the middle insets. Right: Vortex defects also show up as distinct minima in the local energy density on the lattice. We show coarse-grained plots to minimise the influence of lattice artefacts. Lengths are shown in units of $Q^{-1}$, cf. Eq. (3). The healing length $\xi_{\mathrm{h}}=p_{\xi_{\text{h}}}^{-1}=[2g\rho_{0}M]^{-1/2}$, which gives the scale of the diameter of the vortex cores, is $\xi_{\text{h}}\approx 0.83\,Q^{-1}$.
@@ III.1 Low-energy effective theories of scalar fields
Relativistic $\mathrm{O}(N)$ scalar field theories are known to display nonrelativistic effects at low momenta in the presence of a mass gap. For the initial conditions considered here, such a mass gap $|p|\lesssim M$ with effective mass $M$ emerges even when the underlying theory is originally massless Piñeiro Orioli et al. (2015), giving rise to nonrelativistic dynamics in the infrared. Hence, a single component field theory $\mathrm{O}(1)$ can be described by an emergent $\mathrm{U}(1)$ theory. This mapping can be made more rigorous by introducing nonrelativistic degrees of freedom emerging from the underlying relativistic ones
@@
with $\omega_{x}=\sqrt{M^{2}-\nabla^{2}}$ Namjoo et al. (2018); Deng et al. (2018). The emergent dynamics is governed by the Gross-Pitaevskii equation,
@@
thus describing the real-time dynamics of a single-component Bose gas. While it is a well-known mapping, Ref. Deng et al. (2018) considered it for the case of overoccupied isolated systems corroborated by numerical simulations in two dimensions, with initial conditions very similar to what is being considered here.
In contrast, for a nonrelativistic $N$-component Bose gas, a low-energy effective theory was considered in Ref. Mikheev et al. (2019), describing interacting Goldstone modes of the total and relative phase excitations. There are $N$ equations of motion for the $\mathrm{U}(N)$ theory, which give rise to N1N-1 Goldstone modes [^1] Based on the underlying $U(N)$ theory, there are 2N12N-1 broken generators, and 2N22N-2 of them combine into the N1N-1 quadratic modes Mikheev et al. (2019). with quadratic dispersion $\omega_{\mathrm{G}}(p)$ (relative phase excitations), and a Bogoliubov mode (total phase) with dispersion $\omega_{\mathrm{B}}(p)$,
@@
where $\rho_{0}$ is the condensate density in the zero-mode. In Ref. Boguslavski and Piñeiro Orioli (2020), these dispersions were generalised to the relativistic case of O($N$) models by computing the dispersions of possible excitations using classical field equations of motion. This gives rise to the gapless modes
@@
where $g\rho_{0}$ imprints the magnitude of the (local) condensate and is related to the effective mass $M$. These dispersion relations are provided as relative excitations on top of the local condensate that is the solution of the classical equation of motion. Therefore, they can be directly compared to the dispersions (16) of the nonrelativistic theories. However, we emphasise that in the relativistic case, one computes correlations of the scalar field $\phi$ instead of $\psi$ and, before comparing to the stated dispersions, one has to remove the rotation frequency $\pm M$ in the frequency spectrum Boguslavski and Piñeiro Orioli (2020) that also emerges in the exponent of Eq. (14). We will take this into account in all the figures below by plotting $\omega-M$.
@@
We note that in the large-$N$ limit, the Bogoliubov mode is suppressed in the infrared. In this limit, large-$N$ kinetic theories Piñeiro Orioli et al. (2015); Walz et al. (2018); Chantesana et al. (2019) that dominantly describe the scattering of quasiparticles with a quadratic dispersion agree with the low-energy effective theory (EFT) considered in Ref. Mikheev et al. (2019). For the single-component scalar field theory, the EFT considers only Bogoliubov modes and no quadratic excitations. However, the extraction of the excitation spectrum in $\mathrm{U}(1)$ and $\mathrm{O}(1)$ theories in Refs. Piñeiro Orioli and Berges (2019); Boguslavski and Piñeiro Orioli (2020) reveals the existence of a dominant “transport peak” or “non-Lorentzian peak” at low momenta, in addition to the Bogoliubov modes, that drives the dynamics. This raises the question of what kind of quasiparticles are present in a single-component scalar field theory and, in particular, what the dominant degree of freedom is that none of the existing EFTs or kinetic theories are able to capture.
@@ III.2 Topological defects in $\mathrm{O}(1)$ scalar field theory
Approaching from a different perspective, an important consideration is the potential existence of topological defects in our systems. This could be especially significant for the $N=1$ case, since the lower $N$ is, the more constrained the topology of the underlying vacuum manifold becomes, making it more likely to support nontrivial topological structures. [^2] For a mathematically more rigorous consideration of what plays the role of a vacuum manifold in the nonequilibrium dynamical theory considered here, see Refs. Moore (2016); Noel and Spitz (2024).
Due to the effective conservation of particle number in the infrared regime, which also occurs in the presence of a mass gap in the relativistic theories considered here, an inverse self-similar particle transport develops. This gradually transfers particles to low-momentum modes, ultimately building a condensate at zero momentum. However, causality imposes a limit on the rate at which information and correlations can propagate, preventing the instantaneous formation of a uniform condensate. Therefore, the condensate initially lacks long-range order, which it will gradually develop through ordering dynamics Bray (1994). This dynamics may include topological defects depending on the number of field components $N$ Moore (2016), and for a single component, $N=1$, there are stable vortex defects. In this particular case, the formation of vortices and antivortices, and their annihilation process is of importance, since this allows for the growth of larger structures.
One way to identify these defects in our simulations would be performing the canonical transformation (14) on the relativistic degrees of freedom $\phi$ and $\pi$ to obtain the complex $\psi(t,\mathbf{x})$. Then positions on the real-space lattice where depressions in the density of $|\psi(t,\mathbf{x})|^{2}$ occur along with a phase change of approximately $\pm 2\pi$ in $\arg[\psi(t,\mathbf{x})]$ would correspond to vortices and antivortices, respectively Deng et al. (2018). However, we find that it is also possible to obtain topological defects directly from the relativistic degrees of freedom $\phi$ and $\pi$ without transforming into $\psi$, and also without applying any cooling effects in the simulations. For a relativistic theory with a sufficiently large mass gap, defects also show up as minima in energy density fluctuations Noel and Spitz (2024), which coincide with a $\pm 2\pi$ winding in the relativistic phase angle defined by
@@
as shown in Fig. 1 for the two-dimensional case. We have cross-checked that the number and positions of (anti)vortices coincide in these two observables, the relativistic $\theta(t,\mathbf{x})$, and the nonrelativistic $\arg[\psi(t,\mathbf{x})]$. Such phase singularities in the nonrelativistic $\mathrm{arg}[\psi(t,\mathbf{x})]$ have been specifically matched with depressions in the complex field $\psi$ in Refs. Karl and Gasenzer (2017); Deng et al. (2018) and their time evolution has been tracked in the context of vortex-antivortex annihilation events and the resulting coarsening of the domains. In this work, we will obtain results from purely relativistic quantities, namely the fields $\phi(t,\mathbf{x})$ and $\pi(t,\mathbf{x})$ entering $\theta(t,\mathbf{x})$.
Technically, one could first identify the minima in the energy density and check for a corresponding phase winding in $\theta(t,\mathbf{x})$. In practice, however, nearly all the $\pm 2\pi$ phase windings coincide with a minimum in the energy density, such that simply checking only one of these observables gives an accurate picture of the total vortex density. Similarly, for the three-dimensional case, vortex lines are revealed by looking for depressions in the energy density fluctuations, as seen in Fig. 2 for different times.
Figure 3: Initial spatial configurations of the fields $\phi(t_{0},\mathbf{r})$ and $\pi(t_{0},\mathbf{r})$ in $d=2$ dimensions, $\mathbf{r}=(x,y)$, normalised to have values between 1-1 and $1$. The “scrambled” pattern, containing many small domains, allows for numerous phase singularities in $\arg[\phi+\mathrm{i}\pi]$, which we identify as (anti)vortex defects. These domains then become larger, as throughout the coarsening process vortices annihilate with antivortices resulting in the growth of structures. The healing length is $\xi_{\text{h}}\approx 0.83\,Q^{-1}$ in the units shown.
Figure 4: Distribution function $f(t,p)$, cf. (6), for a single-component scalar field theory in (a) two and (b) three dimensions, shown at five different evolution times $t$, where quantities are shown in units of $Q$, i.e., time in units of $Q^{-1}$ and momentum in units of $Q$. The inset shows the functions rescaled according to (22), with $\beta=0.25$ and $\alpha=0.5$ in two dimensions, and $\beta=0.5$ and $\alpha=1.5$ in three dimensions. We also show approximate power law fits $p^{-\kappa}$ in both cases. In (a), simulations were run on a $2048^{2}$ lattice, while in (b) the lattice size was $512^{3}$. The momentum scales $p_{\xi_{\text{h}}}=\sqrt{2g\rho_{0}M}$ corresponding to the healing length $\xi_{\mathrm{h}}=p_{\xi_{\text{h}}}^{-1}$ are $p_{\xi_{\text{h}}}\approx 1.2\,Q$ in $d=2$ and $p_{\xi_{\text{h}}}\approx 0.6\,Q$ in $d=3$.
@@
It is also important to note that using a smoother cutoff for the initial momentum space distribution instead of a step function, like a hyperbolic tangent function or a Gaussian function, did not reduce the number of vortex defects initially formed in the system, which we explicitly checked. As long as one has a relatively steep increase in occupation number in momentum space near some characteristic momentum, this initially scrambled picture will emerge in position space, and vortices will always form very quickly in the dynamics. In this sense, it is not possible not to have, or to remove these defect configurations. For the case of $N=1$ (and also $N=2,3$ but not higher $N$), they are stable configurations on topological grounds Moore (2016); Noel and Spitz (2024).
@@ III.3 Excitations on vortices
Another type of contribution yet to be considered is that the vortex defects themselves can support excitations, which appear as quasiparticles during the dynamics. For instance, phonons can be produced from vortex-antivortex annihilation processes Barenghi et al. (2001); Richaud et al. (2023), which are expected to yield a linear dispersion relation at low momenta. Beyond this, individual vortex lines in three dimensions, and vortex cores in two dimensions can be excited in a way that can be characterised as quasiparticles, which are distinct from the phonons arising from vortex annihilation. In anticipation of our findings, we briefly summarise the properties of such kelvon quasiparticles.
In three dimensional systems, vortex defects are extended objects and may be viewed as vortex lines, along which helical excitations can propagate as waves. These are known as Kelvin waves Thomson (Lord Kelvin), and were originally described in the context of classical hydrodynamics. In the realm of quantum fluids, Kelvin waves arise as helical perturbations of quantised vortex lines, with the circulation being discretised as a direct consequence of the quantum mechanical phase winding around the vortex core Pitaevskii (1961); Roberts (2003a, b); Donnelly (2005). These excitations play a crucial role in the dynamics of quantum turbulence, mediating energy transfer from larger to smaller scales Vinen (2001); Vinen et al. (2003); Kozik and Svistunov (2004, 2005a, 2005b, 2008, 2009); L’vov and Nazarenko (2010); Krstulovic (2012); Laurie et al. (2010); Boué et al. (2011, 2015), as Kelvin waves at a similar scale interact, exciting smaller scale Kelvin waves, ultimately dissipated by phonon emission Vinen et al. (2003); Kozik and Svistunov (2005b); Vinen (2001). At larger scales, vortex reconnection is thought to dominate energy transfer Meichle et al. (2012), and such reconnection events excite Kelvin waves at scales close to the inter-vortex distance. Lord Kelvin obtained, for the small-amplitude wave excitations of a thin columnar vortex with a hollow core, with wave number $p$ along the vortex line and azimuthal quantum number $n=1$, the dispersions Thomson (Lord Kelvin)
@@
where $\gamma\approx 0.577216\dots$ is the Euler-Mascheroni constant. The Kelvin-wave dispersion (21) also applies to vortices in a superfluid Donnelly (2005) with integer winding number ${w}\in\mathbb{Z}$ and circulation $\Gamma={w}h/m$ quantised in integer multiples of the Planck constant $h$ over particle mass. The much larger $\omega^{+}(p\to 0)\simeq\Gamma/(\pi r_{\mathrm{c}}^{2})$ is related to the Magnus force and the mass inside the core Giuriato et al. (2020); Richaud et al. (2021).
@@
The dynamics of quantised vortices and their Kelvin-wave excitations with dispersion (20) is well captured by the Gross-Pitaevskii equation (GPE) (15) Simula et al. (2008a, b); Simula (2013); Baggaley and Laurie (2014); Clark Di Leoni et al. (2015). As it is explicitly visible in the approximate expression (21), the $p$-dependence of the dispersion (20) is nearly quadratic at low momenta. Around $pr_{\mathrm{c}}=1$ it is close to linear in $p$, while at higher momenta, it bends to a lower power than linear. Such a momentum dependence serves as a unique signature of Kelvin waves that is unlikely to be attributed to any other excitation with only linear or quadratic dispersions. Previous GPE studies of Kelvin waves mostly focused on weak wave turbulence, and a direct connection has not been made with nonthermal fixed points to date, where strong turbulence, including vortex tangles, dominates the infrared region.
@@
At first sight, Kelvin waves cannot occur in two dimensions, where vortices lack a linear extension along the core and are localised objects with circular symmetry. Similar quasiparticles, however, emerge even in two dimensions Simula (2018, 2020), especially if the vortices are not point-like, but have a finite core size, which is precisely our case, as seen in the right panel of Fig. 1 depicting the energy density. As a result, the boundary of the vortex supports bound-state-like oscillations, which, in the azimuthal direction, are quantised in the same way as kelvons in three dimensions, looked at in a single fixed plane perpendicular to the vortex core. The lowest such azimuthal excitation corresponds to a circularly varying displacement of the centre of the vortex. As a result, also the density is modified in its radial dependence relative to the rotating position of the vortex, and in the large-volume limit, a continuum of kelvon states emerges, which can be distinguished again by a wave number $p$ parametrising the radial dependence as well as the dispersion of the kelvon. Therefore, an analogy can be drawn between a Kelvin wave extended along a vortex line in three dimensions and a “Kelvin wave” representing, at lowest order, a circularly rotating displacement of the vortex position. We will also refer to these quasiparticles in two dimensions as kelvons, for which a dispersion with the same scaling (21) as in three dimensions has been found Simula (2018, 2020). Wave-like motion along vortex boundaries has also been discussed in the context of Kelvin waves in finite-size, quasi two-dimensional systems as realised in a trapping potential strongly confined in the third dimension Simula (2018, 2013).